OpenStudy (anonymous):

In traveling from city A to city B, Jojo drove for 1 hour at 50 kph and for 3 hours at 60 kph. What was his average speed for the whole trip? please help! thanks :)

OpenStudy (anonymous):

(50X1+60X3)/(1+3) =230/4=57.5

OpenStudy (ajprincess):

Average speed=Total distance/total time Distance=Time*speed

OpenStudy (anonymous):

Average Speed = Total Distance / Total time taken To find total distance covered, lets use this equation: d = v x t d = 50 x 1 = 50m **** This is the distance covered when he was travelling for 1 hour + d = 60 x 3 = 180m **** This is the distance covered when he was travelling for 3 hours So lets add the distances together 50 + 180 = 230m Great! So lets use our original equation for finding the average speed. Average speed = 230 / 4 = 57.5 kmph :)

OpenStudy (anonymous):

thanks for your help :)